Subtract the complex expressions, and simplify.
(-6.2 + 5.9i) – (2.4 + 9.1i)

Answer by Tatiana_Stebko(1539)

(-6.2 + 5.9i)–(2.4 + 9.1i)= -6.2 +5.9i – 2.4 - 9.1i=-8.6-3.2i